### v2

# Mapping tuberculosis fatalities by region and age group in South Korea: A dataset for targeted health policy optimization

Abstract
In South Korea, age-disaggregated tuberculosis (TB) data at the district level are not publicly available due to privacy constraints, limiting fine-scale analyses of healthcare accessibility. To address this limitation, we present a high-resolution, district-level dataset on tuberculosis fatalities and hospital accessibility in South Korea, covering the years 2014 to 2022 across 228 districts. The dataset is constructed using a reconstruction method that infers age-disaggregated TB cases and fatalities at the district level by integrating province-level age-specific statistics with district-level spatial and demographic data, enabling analyses that account for both spatial heterogeneity and age structure. The reconstructed dataset supports temporal analyses of TB burden, hospital availability, and demographic variation over time, and provides a resource for spatial epidemiology and healthcare accessibility studies that require both fine spatial resolution and demographic detail.
Background & Summary
Tuberculosis (TB) remains one of the leading causes of infectious disease mortality worldwide, with an estimated 1.3 million deaths annually [1, 2]. Despite long-term global efforts to control the disease, substantial regional disparities in TB incidence and outcomes persist [3, 4]. Within the Organization for Economic Co-operation and Development (OECD), South Korea presents a notable case: it consistently reports over 10,000 new TB cases annually and ranks second in TB incidence and fifth in TB-related mortality among OECD countries [5]. This combination of persistently high burden and comprehensive public health surveillance distinguishes Korea from most high-income countries with small case numbers.
South Korea maintains a fully digitized public health reporting system in which annual statistics on TB patients, fatalities, and healthcare infrastructure are collected and released at the administrative district level [6, 7], covering more than 200 districts nationwide. In such data-rich settings, the integration of population and geographic information enables detailed analyses of spatial inequality and accessibility in urban systems [3, 4, 8, 9, 10]. More broadly, the spatial distribution of public infrastructure has been studied as a generic feature of urban systems, where facility density scales with population density [10, 13, 15, 12, 16]. In the context of healthcare, such spatial organization can influence both accessibility and robustness under uneven demand or localized constraints [9, 17].
Age is a critical determinant of TB disease severity [5, 18]. Older individuals are more likely to experience fatal outcomes and may face structural barriers to early diagnosis or continuous treatment [19, 20]. However, due to privacy regulations, Korean public health statistics do not provide TB data that are jointly disaggregated by age and higher-spatial-resolution administrative districts [6]. While province-level age disaggregation is available at lower spatial resolution, this limitation creates a structural data gap that hinders age-aware spatial analysis and equity assessments at the district level.
To address this gap, we reconstruct a district-level, age-disaggregated dataset of TB patients and fatalities spanning 2014 to 2022. Our method integrates publicly available province-level age distributions with district-level totals through an upscaling procedure, thereby introducing age resolution while preserving spatial fidelity. Because the reconstruction is based solely on aggregated statistics and does not involve individual-level records, it does not increase privacy risks beyond those present in the original data. The full reconstruction pipeline and source code are released alongside the dataset, enabling reproducibility and reuse in spatial epidemiology and healthcare accessibility studies.
Figure 1 summarizes the demographic motivation for the dataset. The TB fatality rate, defined as the number of TB deaths divided by the number of TB patients, increases strongly with age [Fig. 1(a)]. Since the average age of the general population varies substantially across districts [Fig. 1(b,c)], TB fatality outcomes are expected to differ across regions when age structure is taken into account. These observations motivate analyses of TB burden and healthcare accessibility using data that jointly capture age structure and spatial resolution.
Figure 1: (a) Quartile plots of the TB fatality rate $\phi_{t}$ by age group $t$ across 16 provinces in 2022 (excluding Sejong-si due to the absence of reported fatalities). (b) Average age of the general population across districts. Metropolitan districts tend to have younger populations, while non-metropolitan districts are characterized by older populations. (c) Distribution of the district-level average age, showing demographic heterogeneity across South Korea.
Methods
This section describes the data sources and reconstruction procedure used to generate the dataset. We describe how the district-level age-disaggregated TB data are constructed from the raw statistics provided by KDCA and KOSIS. The overall procedure is summarized in Fig. 2.
Figure 2: Flowchart of the data construction process. Starting from the raw source data (green), we reconstruct the district-level age-disaggregated quantities (pink). By construction, the reconstructed and raw quantities satisfy the relations $N_{s}=\sum_{t}N_{s,t}$ and $D_{s}=\sum_{t}D_{s,t}$. The dataset up to this stage corresponds to the records summarized in Table 1. For subsequent analyses, the dataset is further filtered using empirical criteria. The final dataset (yellow) contains districts $s$ that satisfy $H_{s}\geq 1$ and $N_{s,t},D_{s,t}\geq 1$ for all age groups $t$ of 40 years and older, corresponding to Table 2.
Raw data sources and spatial resolution
We utilize publicly available datasets reported annually from 2014 to 2022, provided by KDCA [6] and KOSIS [7]. Over this nine-year period, data have been collected for 17 provinces and 228 districts. Demographic information, such as the number $N$ of newly reported TB cases and the number $D$ of TB-related deaths by age group, is available at the provincial level, whereas non-demographic information, including the number $H$ of hospitals, is available at the district level. As of 2022, South Korea comprises 228 “si-gun-gu” districts (municipal government level, higher-resolution administrative units), which belong to 17 “si-do” provinces (regional local government level, lower-resolution units). Each province includes as few as 2 districts (Jeju Special Self-Governing Province) and as many as 31 districts (Gyeonggi-do, part of the Seoul metropolitan area). Here, “hospital” refers to a secondary care hospital equipped to provide appropriate treatment for TB. The annual totals of these source datasets are summarized in Table 1. The province-level TB counts and district-level hospital counts form the basis for the reconstructed dataset.
Year NN $D$ $H$
2014 34 869 2 136 330
2015 32 182 2 018 335
2016 30 892 2 020 339
2017 28 161 1 678 344
2018 26 433 1 657 346
2019 23 821 1 492 354
2020 19 933 1 222 360
2021 18 335 1 324 364
2022 16 264 1 223 374
Table 1: Total annual records of the TB source data. Age-disaggregated counts of newly reported TB cases (NN) and TB-related deaths ($D$), covering all age groups, are available only at the province level due to privacy concerns. The number $H$ of hospitals is provided at the district level. Here “hospital” refers to a secondary care hospital equipped to provide proper treatment for TB.
Age distribution statistics
To examine the age-group composition of TB cases and fatalities, we measure the fractions $n_{i,t}$ and $d_{i,t}$ of individuals in age group $t$ among newly reported TB cases and TB-related deaths in province $i$, respectively, as
$n_{i,t}={N_{i,t}}/{N_{i}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ d_{i,t}={D_{i,t}}/{D_{i}},$ (1)
where $N_{i}=\sum_{t}N_{i,t}$ and $D_{i}=\sum_{t}D_{i,t}$. Ages are categorized into ten-year groupings. Similarly, the quantities $N_{t}$ and $D_{t}$, used to compute the age-dependent fatality rate $\phi_{t}$ shown in Fig. 1(a), are calculated as $N_{t}=\sum_{i}N_{i,t}$ and $D_{t}=\sum_{i}D_{i,t}$, respectively.
Reconstruction of district-level age-disaggregated data
The raw demographic data related to TB are officially available in an age-disaggregated format only at the provincial level. To enable a more fine-grained analysis while maintaining statistical robustness, we reconstruct TB case and fatality counts at the district level. For a district $s$ belonging to province $i$, we assume that the age-group fractions are homogeneous within the province, such that $n_{s,t}=n_{i,t}$. Under this assumption, the numbers of TB cases and deaths in district $s\in i$ for age group $t$ are estimated as
$N_{s,t}\equiv n_{i,t}N_{s},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ D_{s,t}\equiv d_{i,t}D_{s},$ (2)
where $N_{s}$ and $D_{s}$ denote the age-aggregated numbers of newly reported TB cases and TB-related deaths in district $s$, respectively, that are available in KDCA and KOSIS. The reconstructed data satisfy $\sum^{\prime}_{s}N_{s,t}=N_{i,t}$ and $\sum^{\prime}_{s}D_{s,t}=D_{i,t}$, where the primed summation runs over districts in province $i$.
Data filtering and inclusion criteria
As shown in Fig. 1(a), we present TB fatality rates only for age groups aged 40 and above, since the rates for younger groups are negligible. Given the steep age dependence observed in the figure, it is also reasonable to expect that fatality rates below age 40 are negligible. Therefore, we focus on five age groups, $t\in\{1,2,3,4,5\}$ (with $t=1$ denoting the age group “40–49”). In addition, district-year records are retained only when they satisfy the conditions of having at least one newly reported TB case ($N\geq 1$), at least one TB-related death among individuals aged 40 years and older ($D\geq 1$), and at least one hospital ($H\geq 1$). The resulting annual summary of reconstructed district-level data is presented in Table 2.
Year $S$ NN $D$ $H$
2014 143 32 322 1 718 328
2015 137 21 683 1 296 312
2016 132 20 207 1 255 306
2017 143 19 427 1 110 330
2018 142 18 145 1 157 334
2019 119 14 318 901 285
2020 116 12 079 736 276
2021 113 10 900 749 262
2022 124 11 789 886 314
Table 2: Summary of annual TB data at the district level available from 2014 to 2022. For each year, the table lists the number $S$ of districts considered, the number $N$ of newly reported TB cases across those districts, the number $D$ of TB-related deaths, and the number $H$ of hospitals. The numbers of newly reported cases $N$ and fatalities $D$ are estimated in an age-disaggregated manner based on province-level age distributions using Eq. (2). All included districts satisfy the conditions of having at least one newly reported TB case ($N\geq 1$), at least one TB-related death among individuals aged 40 years and older ($D\geq 1$), and at least one hospital ($H\geq 1$).
Data Records
The dataset supporting this study has been deposited in the Dryad Digital Repository and is available for peer review at the private link provided below. The released dataset contains annual TB statistics across administrative districts in South Korea from 2014 to 2022. It includes district-level counts of hospitals, TB patients, and fatalities, together with age-stratified proportions of new cases and deaths reported at the provincial level and the reconstructed district-level age-disaggregated quantities obtained from Eq. (2).
At the source-data level, the dataset covers 17 provinces and 228 districts (Table 1). The Dryad repository contains three province-level source files and two sets of district-level files. The province-level files provide annual TB fatalities, newly reported TB cases, and hospital counts for 17 provinces from 2014 to 2022, corresponding to Table 1. Province-level source files provide counts of fatalities and newly reported cases by 5-year age groups, whereas hospital counts are provided without age stratification.
At the district level, the repository provides one full reconstructed file for each year (sigungu_nd_age_YYYY.csv, 2014–2022) and one filtered analysis file for each year (sigungu_nd_age_filtered_YYYY.csv, 2014–2022). The full district-level files (Table 1) reproduce the province-level counts, except for Sejong-si, when aggregated to the provincial level. It contain district names, province names, hospital counts, district area, total newly reported TB cases, total TB-related fatalities, and reconstructed age-group proportions for new cases and fatalities. The filtered files (Table 2) are derived from the full district-level files and retain only districts with at least one newly reported TB case, at least one TB-related death, and at least one hospital, while restricting the age-resolved variables to age groups 40 years and older. At the reconstructed district-year level, the number of included districts ranges from 113 to 143 per year after applying the inclusion criteria summarized in Table 2.
In the full district-level files, the variables n0_9 to n80_ denote the age-group proportions of newly reported TB cases, and d0_9 to d80_ denote the age-group proportions of TB-related fatalities in Eq. (1), for 10-year age bins from 0–9 to 80+. The filtered district-level files include the corresponding variables only for age groups 40–49 through 80+, consistent with the analytical scope of the present study. District area is reported in square kilometers, and hospital variables correspond to the total number of hospitals recorded for each administrative district and year. A complete description of file contents, variable names, and age-group definitions is provided in the accompanying README and repository metadata, allowing users to distinguish source-level provincial counts from reconstructed district-level age-resolved quantities.
Technical Validation
Figure 3: Temporal patterns of the source data for (a) the number $N$ of newly reported TB cases, (b) the number $D$ of TB-related deaths, and (c) the number $H$ of hospitals. The 16 provinces with sufficient data are represented by distinct symbols.
Figure 4: Regional and age-grouped tendency of (a) TB patient fraction $n$, (b) TB death fraction $d$, and, for comparison, (c) the general population demographic fraction $g$ for each province in 2022. The 16 provinces are represented by distinct symbols.
Temporal consistency of source data
Figure 3 shows the temporal patterns of the source data for each province separately. Seoul-si and Gyeonggi-do exhibit markedly higher values across all measured quantities than other provinces, reflecting their large populations. The number of newly reported TB cases $N$ shows a consistent downward trend across provinces [Fig. 3(a)], and the number of TB-related deaths $D$ also decreases with small fluctuations [Fig. 3(b)]. The number of hospitals $H$ remains nearly unchanged in most provinces, except for Gyeonggi-do, where hospital counts increase modestly during the study period. These temporal patterns are preserved in the released dataset.
Age distribution verification
The age-group fractions defined in Eq. (1) are shown in Fig. 4. The TB-related fractions $n$ and $d$ increase monotonically with age, with a stronger age dependence for TB-related deaths than for newly reported cases [Fig. 4(a,b)]. For comparison, the age distribution of the general population is shown in Fig. 4(c). Although the population is concentrated in the 40–60 year age range, TB-related deaths (cases) are disproportionately concentrated among individuals aged 70 years and older. This comparison indicates that the increasing age pattern of TB-related quantities is not simply a reflection of the underlying population structure. Instead, it reveals a distinct age pattern specific to TB outcomes. These observations motivate the use of TB-related age fractions, rather than the general population age distribution, when constructing the age-disaggregated dataset. These age-specific patterns are also consistent with the decision to focus the filtered analysis dataset on individuals aged 40 years and older, for whom TB-related fatality is non-negligible.
Consistency of reconstructed counts
The reconstructed district-level data satisfy the aggregation constraints in Eq. (2) by construction. Summing $N_{s,t}$ and $D_{s,t}$ across districts within each province reproduces the province-level totals $N_{i,t}$ and $D_{i,t}$ reported in the original statistics. This ensures that the reconstructed district-level age-resolved dataset remains fully consistent with the officially reported province-level counts. In addition to the aggregate consistency described above, the released data files were organized so that the provenance of each variable can be traced directly to the underlying source or reconstruction step. Province-level files contain the original annual TB statistics by province, while district-level files store the reconstructed age-resolved quantities together with district-level totals of hospitals, TB cases, fatalities, and area. The filtered district-level files were generated deterministically from the full district-level files by applying the same inclusion criteria used in this paper.
Usage Notes
The reconstructed dataset is intended for comparative and exploratory studies of TB burden, healthcare accessibility, and demographic heterogeneity across South Korea. Because age-disaggregated district-level counts are inferred from province-level age fractions, the reconstructed quantities should be interpreted as estimates under the assumption of within-province homogeneity in age composition. Hospital accessibility in the dataset is represented through district-level hospital counts and areal hospital density. Such quantities provide a useful proxy for infrastructure availability, but they do not explicitly account for regional differences in hospital capacity, quality of care, or transportation infrastructure. The dataset may therefore be combined with other mobility or infrastructure data in future studies. In addition, the dataset can be used together with previously developed spatial allocation frameworks that analyze district-level hospital accessibility and TB fatalities [11].
Potential research application example: Healthcare facility optimization
As an illustrative application, the dataset can support studies on the spatial optimization of healthcare facilities. For example, the dataset can be combined with the hospital allocation framework proposed in Ref. [11]. To optimize hospital density for minimizing TB fatalities, they formulated the TB fatalities $E$ as a function of hospital density while accounting for human mobility, based on a random-walk model with traps [21], as follows:
$E({\vec{\eta}})=\sum_{s}{N_{s}\phi_{s}(\eta_{s})=\sum_{s}{N_{s}\exp(-\eta_{s}/\tilde{\eta}_{s})}},$ (3)
where the hospital density $\eta_{s}$ and the fatality rate $\phi_{s}$ for district $s$ are defined as $\eta_{s}\equiv H_{s}/A_{s}$ and $\phi_{s}\equiv D_{s}/N_{s}$, where $H_{s}$ is the number of hospitals, $A_{s}$ is the district area, and $D_{s}$ and $N_{s}$ are the numbers of TB-related deaths and newly reported TB cases, respectively, in a given year. The hospital configuration across all districts is denoted by $\vec{\eta}=(\eta_{1},\eta_{2},\cdots,\eta_{s},\cdots)$. The characteristic hospital density $\tilde{\eta}_{s}$, which reflects the medical and infrastructural environment, can be empirically determined by equating the observed fatality rate ($\phi_{s}\equiv D_{s}/N_{s}$) to the modeled form $\phi_{s}=\exp(-\eta_{s}/\tilde{\eta}_{s})$, yielding $\tilde{\eta}_{s}=\eta_{s}/\log(N_{s}/D_{s})$, as reported in the previous study [11]. Preserving the total number of hospitals, the optimal hospital density of district $s$ is theoretically obtained as $\eta_{s}^{\rm(opt)}=\tilde{\eta}_{s}\log\left({\rho_{s}}/{(z\tilde{\eta}_{s})}\right)$, where $\rho_{s}={N_{s}}/{A_{s}}$ denotes the patient density and $z$ is a Lagrange multiplier. The previous study further demonstrated that the ratio ${\eta^{\rm(opt)}}/{\eta}$ between the optimal and current hospital densities strongly depends on the characteristic patient density ${\rho_{s}}/{\tilde{\eta}_{s}}$ [11]. This modeling framework assumes that the fatality rate $\phi$ depends solely on the characteristics of each district and therefore does not capture the well-known dependence of TB fatality on patient age [1, 18]. This simplification is primarily due to data limitations. While TB data at the district level are more suitable for statistically robust analysis, age-disaggregated TB data are only available at the province level due to privacy concerns.
Figure 5: (a) Illustrative example of how the released dataset can be used in spatial healthcare allocation models. District-level values of the difference ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$ across South Korea are visualized. The inset shows a magnified view of Seoul. Uncolored districts indicate no data satisfying the criteria. (b-c) Comparison of the difference ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$ for $\rho_{s,5}$ (b) and $\rho_{s,5}/\tilde{\eta}_{s,5}$ (c). No distinct pattern observed along $\rho_{s,5}$, while a steady, roughly linear increase is seen along $\rho_{s,5}/\tilde{\eta}_{s,5}$ despite the log-scaled x-axis. This figure is provided solely as an illustrative example of dataset usage.
Using age-resolved district-level TB data in Eq. (2), it is possible to minimize the total number of TB-related deaths by reallocating hospitals across districts while accounting for age-specific effects. Reformulating the objective function in Eq. (3) by decomposing total fatalities into age-group components, the original fatality rate $\phi_{s}$ is extended to $\phi_{s,t}\equiv D_{s,t}/N_{s,t}$, representing the fatality rate for age group $t$ in district $s$. As stated in Eq. (3), the fatality rate was modeled as $\phi_{s}(\eta_{s})=\exp(-\eta_{s}/\tilde{\eta}_{s})$, where $\eta_{s}$ denotes the areal density of hospitals and $\tilde{\eta}_{s}$ represents a district-specific characteristic density. In the age-disaggregated formulation, individuals in district $s$ access a shared set of $H_{s}$ hospitals regardless of age, implying that the hospital density remains $\eta_{s,t}=\eta_{s}$. However, hospital accessibility may vary across age groups (e.g., due to differences in mobility and activity), resulting in an age-dependent characteristic density $\tilde{\eta}_{s,t}$. Accordingly, the age-specific fatality rate is modeled as
$\phi_{s,t}=\exp\left(-{\eta_{s}}/{\tilde{\eta}_{s,t}}\right),$ (4)
with the characteristic density empirically estimated as
$\tilde{\eta}_{s,t}=\frac{\eta_{s}}{-\log\phi_{s,t}}=\frac{H_{s}/A_{s}}{\log(N_{s,t}/D_{s,t})},$ (5)
using the observed values of $N_{s,t}$, $D_{s,t}$, and $H_{s}$ summarized in Table 2. This formulation implies that $\tilde{\eta}_{s,t}$ is directly proportional to the hospital density and inversely related to the observed fatality rate. For elderly age groups in districts with insufficient hospital infrastructure, $\tilde{\eta}_{s,t}$ tends to be high, implying that more hospitals are required to reduce fatality rates. Consequently, a comparable reduction in $\phi_{s,t}$ requires a larger increase in hospital density than in districts with better infrastructure and access.
Using the model defined in Eq. (4), the total number of fatalities, which serves as the objective function to be minimized, can be written as a function of hospital densities:
$E_{\rm fatalities}(\vec{\eta})=\sum_{s,t}D_{s,t}=\sum_{s,t}N_{s,t}\phi_{s,t}=\sum_{s,t}N_{s,t}\exp\left(-{\eta_{s}}/{\tilde{\eta}_{s,t}}\right).$ (6)
$\vec{\eta}$ is varied to minimize $E_{\rm fatalities}(\vec{\eta})$ while the total number $H$ of hospitals is preserved. The optimal hospital density $\eta_{s}^{\rm(opt)}$ that minimizes $E_{\rm fatalities}(\vec{\eta})$ can, in principle, be obtained analytically using the Lagrange multiplier method. However, when age groups $t$ are taken into account in Eq. (6), a closed-form solution is not available. Therefore, a numerical optimization could be performed to determine the optimized hospital configuration vector $\vec{\eta}^{\rm(opt)}=\left(\eta_{1}^{\rm(opt)},\eta_{2}^{\rm(opt)},\cdots,\eta_{s}^{\rm(opt)},\cdots\right)$ using a zero-temperature Monte Carlo approach. A randomly selected pair $(u,v)$ of districts attempts to exchange a small amount of hospital allocation, denoted by $\Delta H$, through the updates $H_{u}\to H_{u}-\Delta H$ and $H_{v}\to H_{v}+\Delta H$. If this relocation results in a decrease in the total fatalities $E_{\rm fatalities}$, the new configuration is accepted. This process is repeated until $E_{\rm fatalities}$ converges to a minimum value, denoted by $E^{\rm(min)}_{\rm fatalities}$. The optimized density configuration is defined as the one that minimizes the objective function, i.e.,
$E^{\rm(min)}_{\rm fatalities}\equiv E\left(\vec{\eta}^{\rm(opt)}\right).$ (7)
Here $\Delta H=0.01$, which is sufficiently small to ensure the convergence of $E^{\rm(min)}_{\rm fatalities}$ with respect to $\Delta H$.
To visualize the effect of accounting for age-group distributions on the optimization process, the age-considered optimization results are compared to the baseline case studied in the previous work [11], which uses the objective function in Eq. (3). One can first examine the relative change in hospital density compared to the original empirical density $\eta$, specifically ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ and ${\eta^{\rm(opt)}_{\times}}/{\eta}$, which are obtained by minimizing the objective function $E_{\rm fatalities}$ in Eq. (6) for $\eta^{\rm(opt)}_{\rm age}$ and $E(\vec{\eta})$ in Eq. (3) for $\eta^{\rm(opt)}_{\times}$, respectively. Here, the subscript $\times$ represents the optimal hospital density obtained without accounting for age-group differences. Since elderly individuals account for a substantial portion of TB patients and fatalities, one can analyze the optimization results with respect to the areal patient density of the oldest age group ($t=5$, ages 80 and above), defined as $\rho_{s,5}=N_{s,5}/A_{s}$. The relative changes ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ and ${\eta^{\rm(opt)}_{\times}}/{\eta}$ are nearly identical [see the color map in Fig. 5(a) and the y-axis scales in Figs. 5(b) and 5(c)].
A higher density of elderly patients alone does not necessarily lead to a greater hospital allocation in the age-considered optimization [Fig. 5(b)], indicating that absolute patient density is not a sufficient predictor of post-optimization hospital gain. In line with Ref. [11], which showed that the characteristic patient density governs hospital gain or loss, the optimization outcomes from the perspective of the rescaled density $\rho_{s,5}/\tilde{\eta}_{s,5}$ are examined. We find that the rescaled patient density plays a dual role. First, the relative change in hospital density ${\eta^{\rm(opt)}_{\rm age}}/{\eta}$ exhibits a clear dependence on ${\rho_{s,5}}/{\tilde{\eta}_{s,5}}$, indicating that this quantity continues to govern whether a district gains or loses hospitals after optimization [Fig. S1 in the Supplemental Material]. Second, the difference between the age-considered and age-agnostic optimizations, ${(\eta^{\rm(opt)}_{\rm age}-\eta^{\rm(opt)}_{\times})}/{\eta}$, is also systematically controlled by the same rescaled density [Fig. 5(c)]. In this sense, the rescaled patient density ${\rho_{s,5}}/{\tilde{\eta}_{s,5}}$ governs not only the direction of hospital redistribution after optimization, but also amplifies the difference in hospital gains between the age-considered and age-agnostic cases. The characteristic scale $\tilde{\eta}_{s,t}$ defined in Eq. (5) is strongly correlated with the age-aggregated scale $\tilde{\eta}_{s}$, since the empirical fatality rate $\phi_{s,t}$ is constrained by the overall rate $\phi_{s}$. This correlation explains why the behavior of the hospital density difference also mirrors a similar dependence on the age-aggregated ratio ${\rho}/{\tilde{\eta}}$ [see Fig. S2 in the Supplemental Material].
Data Availability
The dataset supporting this study has been deposited in the Dryad Digital Repository and is available for peer review at the following private link: https://datadryad.org/share/qfxSkgpvSxiVsZhnUP8jLxwMMIXUiLwCaVrP5Y0Djek. This dataset contains annual tuberculosis (TB) statistics across administrative districts in South Korea from 2014 to 2022. It includes district-level counts of hospitals, TB patients, and fatalities, along with age-stratified proportions of new cases and deaths (by 10-year age groups from 0–9 to 80+). The dataset was reconstructed using publicly available provincial-level data from the Korea Disease Control and Prevention Agency (KDCA) and the Korean Statistical Information Service (KOSIS), and harmonized with higher-resolution regional information to produce a novel age-stratified TB dataset. All variables are fully documented in the accompanying metadata and README file. The dataset will be assigned a DOI and made publicly available upon publication.
Code Availability
The code used for data preprocessing, age-stratified data reconstruction, and figure generation is openly available at a github repository: https://github.com/kwyosu7/Tuberculosis_hospital_distribution_optimization. The repository includes Python scripts for constructing the dataset and generating the figures used in the manuscript.
Acknowledgements
This work was supported by the National Research Foundation (NRF) of Korea through Grant Numbers. NRF-2023R1A2C1007523 (S.-W.S.), RS-2024-00341317 (M.J.L.), and by KIAS Individual Grants (No. CG079902 (D.-S.L)). We thank APCTP, Pohang, Korea, for their hospitality during the Topical Research Program [APCTP-2025-T04], from which this work benefited greatly.
Author contributions statement
Y.K., D.-S.L., M.J.L., and S.-W.S. designed, conceptualized, and wrote the manuscript. Y.K. and M.J.L. searched out and digitized the TB dataset. D.-S.L. assisted with the early generation of the datasets. Y.K. developed and implemented the optimization models, conducted the computational analysis, and also performed the spatial data processing and visualization. D.-S.L., M.J.L., and S.-W.S. supervised the project and provided critical guidance on the methodological framework. All the authors contributed to reviewing and editing the manuscript.
Competing interests
The authors declare no competing interests.
Supplemental Material
Figure S1: Plot of the ratio $\frac{\eta^{\rm(opt)}_{\rm age}}{\eta}$ versus the rescaled patient density $\frac{\rho_{s,t}}{\tilde{\eta}_{s,t}}$ from 2014 to 2022 (a-i). All age groups exhibit an increasing trend along the log-scaled x-axis.
Figure S2: The difference between relative changes in hospital density $\frac{\eta^{\mathrm{(opt)}}_{\mathrm{age}}-\eta^{\mathrm{(opt)}}_{\times}}{\eta}$ versus the rescaled patient density $\frac{\rho_{s,t}}{\tilde{\eta}_{s,t}}$ from 2014 to 2022 (a-i) in $t=5$ age group and no-age case.
Figure S3: The minimum fatalities $E^{\rm(min)}_{\rm{fatalities}}$ as a function of the age weighting parameter $a$ from 2014 to 2022 is shown representatively, with $a_{\rm min}$ and $a_{\rm max}$. The pink square, green hexagon and purple circle represent the cases of $a_{\rm min}, a=a$$a_{\rm max}$ and $a=0$, respectively.
Figure S4: (a) Relationship between relative changes in optimal hospital allocation $\frac{\eta_{s}^{\rm(opt)}(a\simeq a_{\rm max})}{\eta_{s}}-\frac{\eta_{s}^{\rm(opt)}(a=0)}{\eta_{s}}$ versus fatality $\frac{E_{s}^{\rm(opt)}(a\simeq a_{\rm max})}{D}-\frac{E_{s}^{\rm(opt)}(a=0)}{D_{s}}$ across districts under the $a\simeq a_{\mathrm{max}}$ case compared to the baseline case ($a=0$). (b) A distribution of fatality change $\frac{E_{s}^{\rm(opt)}(a\simeq a_{\rm max})}{D}-\frac{E_{s}^{\rm(opt)}(a=0)}{D_{s}}$ for age group $t$. $\frac{E_{s}^{\rm(opt)}(a\simeq a_{\rm max})}{D}-\frac{E_{s}^{\rm(opt)}(a=0)}{D_{s}}$ for $t=5$ (ages 80 and above) concentrated in the negative range, unlike other age groups.